A rock is shot with a sling shot at angles 30, 45, 60 and 90.

• which launch angle will the rock's hang time be the greatest

• which launch angle will the rock's horizontal and vertical displacement will be the greatest
Answer #1

Initial horizontal velocity of rock ux = u cos θ , vertical velocity of rock u_y=u\sin\theta

Using equation v=u+at

Final horizontal velocity when the rock reaches maximum height is v_x=u_x=u\sin\theta ,final vertital velocity at maximum height is v_y=0

Time taken to reach maximum height is t' .

v_y=u_y-gt'

  \Rightarrow t'=\frac{u_y}{g}=\frac{u\sin\theta}{g}

Time taken to reach the ground is double the time taken to reach maximum height ,t=2 t'=\frac{2u\sin\theta}{g}

For the given angles, \sin\theta will be maximum when \theta=90\degree

Hence rock's hang time will be greatest when \theta=90\degree

------------------------------

Using s=ut+\frac{1}{2}at^2

Rock's maximum vertical displacement y=u\sin\theta-\frac{1}{2}gt'^2 ( time taken to reach maximum vertical displacement is t' )

y=u\sin\theta*\frac{u\sin\theta}{g}-\frac{1}{2}g\left ( \frac{u\sin\theta}{g} \right )^2=\frac{u^2\sin^2\theta}{2g}

for maximum vertical displacement (y) , \sin^2\theta=1 \Rightarrow \theta=90\degree

Rock's maximum horizontal displacement is ( t is total time taken to reach ground)

x=u\cos\theta t=u\cos\theta*\frac{2u\sin\theta}{g}=\frac{u^2 \sin2\theta}{g}

For maximum horzontal displacement (x)  ,  \sin2\theta=1  \Rightarrow \theta=45\degree